two children wrestle over a toy of mass 1.5 kg. The boy pulls with a. Force of 6.0 N(W) while the girl pulls with a force of 8.0 N(E). The toy slides with an acceleration of 1.0m/s2. Determine the value of the frictional force acting in the toy


1
Expert's answer
2021-03-09T15:27:05-0500

"8-6-F_f=ma\to"


"F_f=8-6-ma=8-6-1.5\cdot1=0.5\ (N)" . Answer